# API Pack: A Massive Multi-Programming Language Dataset for API Call Generation

### v3

@@ 4.2 Evaluations
5 Results
Table 3: Evaluation for models fine-tuned with 20k Python API dataset. A comparison with (non-fine-tuned) GPT-3.5 and GPT 4 performance for level 3 is also included.
@@ 5 Results
Mistral-7b zero-shot 0-shot 17.2 10.9 14.1 11.4 14.3 11.2
@@
CodeLlama-7b zero-shot 0-shot 8.1 6.1 10.0 7.0 11.0 7.8
@@
Llama-2-13b zero-shot 0-shot 9.4 6.2 11.6 9.0 10.9 8.4
@@
CodeLlama-13b zero-shot 0-shot 9.8 6.8 10.8 8.1 12.1 8.5
@@
5.1 Fine-tuned CodeLlama Excels in API call Generation
Table 3 presents the evaluation results for the four models fine-tuned with 20,000 Python instances from API Pack. The fine-tuned CodeLlama-13b model demonstrates superior performance in API call generation for the three-shot retrieved setting, achieving the highest API call accuracy across all evaluation levels. Models fine-tuned with the three-shot template consistently outperform those fine-tuned with the zero-shot template, showing the importance of using the three-shot template to improve the model’s in-context learning abilities. Furthermore, prompting with 3-shot (retre) at testing time leads to substantial improvements compared to 0-shot, a trend observed across all models and levels. This finding highlights the significance of providing relevant examples to improve the models’ accuracy in generating API calls. Notably, the fine-tuned CodeLlama-13b model with API Pack surpasses the performance of GPT-3.5 and GPT-4 models (not fine-tuned) for Level 3, both 0-shot and 3-shot (retre).
Table 4: Performance comparison for different retrieval methods in three-shot prompting.
@@ 5.1 Fine-tuned CodeLlama Excels in API call Generation
5.2 Retrieval Augmentation Improves API Call Generation
Table 4 illustrates the impact of different retrieval methods on three-shot API call generation, highlighting the significance of retrieval augmentation in improving the models’ performance. The results of fine-tuning Mistral-7b and CodeLlama-13b models with 20,000 Python instances from API Pack using the three-shot template demonstrate that three-shot retrieval is the most effective approach.
@@ 5.3 Cross-lingual Generalizations in API Call Performance
The graph shows that generalizing to new programming languages without fine-tuning is challenging in zero-shot scenarios. However, the models perform better than expected in three-shot settings, indicating some in-context learning adaptability. Mixed-language fine-tuning improves performance in both zero-shot and three-shot scenarios, suggesting that even a small amount of fine-tuning data in various languages can improve the model’s effectiveness in cross-lingual API call tasks. These findings demonstrate that cross-lingual API call generation can be achieved by fine-tuning models on a large amount of data in one language plus small amounts of data from other languages, without requiring an extensive dataset for each target programming language. Figure 4 illustrates the models’ adaptability across ten programming languages, with the Mixture Model performing on par with Expert Models in three-shot testing, highlighting their potential in multi-lingual programming applications.
@@ 5.4 Scaling Instruction Dataset Helps Generalization
Figure 5: Scaling instruction dataset on CodeLlama-13b with zero-shot and three-shot retre evaluations. The x-axis represents the size of fine-tuning data from API Pack. The y-axis is Endpoint or API call accuracy.
Fig. 5 clearly demonstrates an upward trend in zero-shot performance that correlates with the fine-tuning dataset size (20k, 40k, 80k, 100k, and 1 million instances from API Pack). This improvement highlights the advantage of larger datasets in providing diverse examples, which is crucial in zero-shot prompting where the model relies solely on its pre-existing knowledge. For three-shot prompting, the graph also shows improvements as the dataset size increases, although less significantly than in the zero-shot case, particular for level 3 cases. This suggests that while additional fine-tuning data is beneficial, model pretraining plays a more critical role in in-context learning ability.
@@ 5.5 Additional Ablation Experiments
To verify API Pack’s quality, we evaluated a fine-tuned CodeLlama-13b model on 20k instances of ToolBench dataset, which only contains APIs from RapidAPI. Results in Table 5 show that using APIs from a single source like in TollBench decreases model performance in both 0-shot and 3-shot settings compared to API Pack, which curates data from four different sources.
@@ 6 Conclusion
In this paper, we introduce API Pack, a massive multi-programming language dataset with over 1 million instruction-API call instances. Using this dataset, we investigate if models can generalize to new languages without extensive multi-language data and whether exposure to diverse APIs during fine-tuning improves generalization. Our results show that cross-lingual API call is achievable by training on a large amount of single-language data supplemented with small amounts of data from other languages. Furthermore, increasing data volume improves generalization. These findings highlight the potential of large-scale, diverse datasets in enabling LLMs to effectively identify appropriate API endpoints and generate corresponding API calls based on natural language instructions.
7 Limitations
API Pack is not designed for multi-API call scenarios, so models fine-tuned with this dataset may struggle with tasks involving multiple, interdependent API calls. This limitation restricts API Pack’s effectiveness in complex, real-world software development tasks that often require multi-API integrations. Future work should focus on creating datasets and models that can effectively handle multi-API call scenarios, enabling the generation of more sophisticated and practical code solutions.

### v4

@@ 4.2 Evaluations
5 Experimental Results
5.1 Fine-tuned CodeLlama Excels in API call Generation
Table 3 presents the evaluation results for the four models fine-tuned with 20,000 Python instances from API Pack. The fine-tuned CodeLlama-13b model demonstrates superior performance in API call generation for the 3-shot retrieved setting, achieving the highest API call accuracy across all evaluation levels. Models fine-tuned with the 3-shot-tuning template consistently outperform those fine-tuned with the 0-shot-tuning template, showing the importance of using the 3-shot template to improve the model’s in-context learning abilities during training. Furthermore, prompting with 3-shot (retre) at testing time leads to substantial improvements compared to 0-shot, a trend observed across all models and levels. This finding highlights the significance of providing relevant examples to improve the models’ accuracy in generating API calls. Notably, the fine-tuned CodeLlama-13b model with API Pack surpasses the performance of GPT-3.5 and GPT-4 models (not fine-tuned) for Level 3, both 0-shot and 3-shot (retre).
Table 3: Evaluation for models fine-tuned with 20k Python API dataset, including a comparison with (not-fine-tuned) GPT-3.5 and GPT 4. Note that 3-shot (retre) is short for 3-shot retrieved.
@@ 5.1 Fine-tuned CodeLlama Excels in API call Generation
Mistral-7b 0-shot 0-shot 17.2 10.9 14.1 11.4 14.3 11.2
@@
CodeLlama-7b 0-shot 0-shot 8.1 6.1 10.0 7.0 11.0 7.8
@@
Llama-2-13b 0-shot 0-shot 9.4 6.2 11.6 9.0 10.9 8.4
@@
CodeLlama-13b 0-shot 0-shot 9.8 6.8 10.8 8.1 12.1 8.5
@@
5.2 Retrieval Augmentation Improves API Call Generation
Table 4 illustrates the impact of different retrieval methods on 3-shot API call generation. The results demonstrate that at inference time, 3-shot (retre) is the most effective approach for fine-tuned Mistral-7b and CodeLlama-13b models compared to 3-shot (rand) and 3-shot (retre & rerank). These models were fine-tuned using 20,000 Python instances with the 3-shot-tuning template.
Table 4: Performance comparison for different retrieval methods in 3-shot prompting. Note that 3-shot (retre & rerank) is short for 3-shot retrieved & re-ranked.
@@ 5.3 Cross-lingual Generalizations in API Call Performance
The graph shows that generalizing to new programming languages without fine-tuning is challenging in 0-shot scenarios. However, the models perform better than expected in 3-shot settings, indicating some in-context learning adaptability. Mixed-language fine-tuning improves performance in both 0-shot and 3-shot scenarios, suggesting that even a small amount of fine-tuning data in various languages can improve the model’s effectiveness in cross-lingual API call tasks. These findings demonstrate that cross-lingual API call generation can be achieved by fine-tuning models on a large amount of data in one language plus small amounts of data from other languages, without requiring an extensive dataset for each target programming language. Figure 3 illustrates the models’ adaptability across ten programming languages, with the Mixture Model performing on par with Expert Models in their perspective programming languages in 3-shot testing, highlighting their potential in multi-lingual programming applications.
@@ 5.4 Scaling Instruction Dataset Helps Generalization
Figure 4: Scaling instruction dataset on CodeLlama-13b with 0-shot and 3-shot retrieval evaluations. The x-axis represents the size of fine-tuning data from API Pack. The y-axis is Endpoint or API call accuracy.
Fig. 4 clearly demonstrates an upward trend in 0-shot performance that correlates with the fine-tuning dataset size (20k, 40k, 80k, 100k, and 1 million instances from API Pack). This improvement highlights the advantage of larger datasets in providing diverse examples, which is crucial in 0-shot prompting where the model relies solely on its pre-existing knowledge. For 3-shot prompting, the graph also shows improvements as the dataset size increases, although less prominent than in the 0-shot case, particularly for level 3 cases. This suggests that while additional fine-tuning data is beneficial, model pretraining plays a more critical role in in-context learning ability.
@@ 5.5 Additional Ablation Experiments
To demonstrate the benefit of API Pack compared to existing API datasets in the literature, we evaluated a fine-tuned CodeLlama-13b model on 20k instances of ToolBench dataset, which only contains APIs from RapidAPI. Results in Table 5 show that using APIs from a single source like in ToolBench significantly decreases model performance in both 0-shot and 3-shot settings compared to API Pack, which curates data from four different sources.
@@ 6 Conclusion
In this paper, we introduced API Pack, a multi-programming language dataset containing over 1 million instruction-API call pairs. Our experiments demonstrated that fine-tuning CodeLlama-13B on 20,000 Python instances from API Pack enables it to outperform GPT-3.5 and GPT-4 in generating unseen API calls. We also showed that cross-lingual API call generation is achievable by primarily training on single-language data augmented with a small amount of data from other languages. Furthermore, we found that increasing the training data volume improves the model’s ability to generalize to unseen API calls. These findings highlight the potential of large-scale, diverse datasets in enabling LLMs to effectively identify appropriate API endpoints and generate corresponding API calls based on natural language instructions.
7 Limitations and Future Work
The first release of API Pack exhibits a set of core limitations. First, the fact API Pack was not designed for multi-API call scenarios, limits its application to single-API development tasks. Second, current instruction-style requires of the API Name. Third, programming language coverage is inherently limited by the library that we used to generate code. Finally, only few OAS files contained license information. We plan to address these limitations in future versions of the dataset, by integrating multi-API call scenarios and instructions that simply require a task description. We also plan to explore other methods to generate API calls so that we can further increase programming language diversity, and to track down complete and accurate license information for all APIs included in API Pack.